OpenStudy (emmynimmy):

What is the slope of the line that passes through the points (1, 2) and (-2, 4)? 2 –2 –2/3 –3/2

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

Let me explain ^_^

OpenStudy (emmynimmy):

Thanks! I have another question. A line has a slope of 2 and passes through the point (4, 0). What are the coordinates of another point that is on the line? (4, 6) (6, 4) (2, 0) (0, 2)

Nnesha (nnesha):

formula to find slope \[\huge\rm \frac{ y_2 -y_1 }{ x_2-x_1 }\] plugin in x's and y's values (x_1 , y_1)(x_2,y_2)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Captain_America_Army Let me explain ^_^ \(\color{blue}{\text{End of Quote}}\) next time explain first and let them find the answer.

OpenStudy (emmynimmy):

can someone help me with the second question?

OpenStudy (anonymous):

Sorry I'm trying to remember a Site that helps in graphing

OpenStudy (emmynimmy):

4/5 5/4 –4/5 –5/4

OpenStudy (anonymous):

Whats the question?

Nnesha (nnesha):

that's why u put girls pic as pfp? anywz

Nnesha (nnesha):

emmy use the slope formula

OpenStudy (emmynimmy):

What is the slope of line CD? This is the question to the picture that is attached with it.

OpenStudy (emmynimmy):

A line has a slope of 2 and passes through the point (4, 0). What are the coordinates of another point that is on the line? (4, 6) (6, 4) (2, 0) (0, 2) And I need answer for this one, its a timed assignment.

Nnesha (nnesha):

we have to find 2nd point first one is (4,0) let's try last one (4,0)(0,2)\[\frac{ 2-0 }{ 0-4}=?\] if you get 2 then (0,2) is the 2nd point if it's not equal to 2 then try ohter 3 points